OpenStudy (anonymous):

Trig Identities Proof: \[\frac{ cosx-sinx }{ cosx+sinx }=\sec2x-\tan2x\]

OpenStudy (anonymous):

Please help! :(

Nnesha (nnesha):

cos = ? sin = ?

OpenStudy (anonymous):

I'm not sure what you mean?

Nnesha (nnesha):

trig identities sin cos tan properties tan = sin /cos sin = ? cos = ?

OpenStudy (anonymous):

sin=y/r cos=x/r

Nnesha (nnesha):

o^_^o \[\rm \sin = \frac{ 1 }{ \csc }\] \[\rm \cos = \frac{ 1 }{ \sec }\]

OpenStudy (anonymous):

ooh...

Nnesha (nnesha):

solve L.H.S replace cos by 1/sec and sin by 1/csc

ganeshie8 (ganeshie8):

that works nicely, there is a shortcut though.. multiply top and bottom by the conjugate of denominator : `cosx - sinx`

ganeshie8 (ganeshie8):

and recall the identity : \[\cos^2x-\sin^2x=\cos(2x)\]

Nnesha (nnesha):

okay so right is that \[\large\rm \sec^2 x - \tan^2x ~~or~ \sec(2x) - \tan (2x)\] square or 2x ?

OpenStudy (anonymous):

2x @Nnesha Sorry, I got logged out suddenly and had some difficulty logging back on.

OpenStudy (theopenstudyowl):

@ineedhelpnowplz